The area of a circle 28 square meters. What is its circummference?
Answer by Alan3354(69443)

The area of a circle 28 square meters. What is its circummference?
------
Area = pi*r^2
r = sqrt(Area/pi)
--------
C = 2 pi*r
C = 2 pi(sqrt(Area/pi))
C = 2 pi(sqrt(28/pi)) = sqrt(4pi^2*28/pi)
= sqrt(112pi)
= ~ 18.758 meters